OpenStudy (anonymous):

Choose an equation in slope-intercept form for the line that passes through (0,7) and is parallel to the y=4x-19. Select one: a. y = 4x + 3 b. y = 4x + 7 c. y= (1/4)x - 7 d. y= (1/4)x + 9

jimthompson5910 (jim_thompson5910):

Hint: any two parallel lines have the same slope, so your answer will be an equation with a slope of 4 (since y = 4x - 19 has a slope of 4)

jimthompson5910 (jim_thompson5910):

you just need to make sure that the equation goes through (0,7)

OpenStudy (anonymous):

thank you! i will try it now.

jimthompson5910 (jim_thompson5910):

ok tell me what you get

OpenStudy (anonymous):

i don't get how to set it up.

jimthompson5910 (jim_thompson5910):

y = mx + b y = 4x + b ... we know the slope is 4, so plug it in

jimthompson5910 (jim_thompson5910):

we also know that it must go through (0,7), so plug in x = 0 and y = 7 y = 4x + b 7 = 4(0) + b 7 = 0 + b 7 = b b =7 So because m = 4 and b = 7 we go from y = mx+b to y = 4x + 7

OpenStudy (anonymous):

i got it :) thank you!

OpenStudy (anonymous):

now i know how to set them up

jimthompson5910 (jim_thompson5910):

that's great, yw